OpenStudy (anonymous):

Write an algebraic expression that is equivalent to sin(tan−1x2).

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

is it \[\sin(\tan^{-1}(x^2))\]?

OpenStudy (anonymous):

Write an algebraic expression that is equivalent to sin(tan^−1(infinity/2)).

OpenStudy (freckles):

you know infinity isn't a number right?

OpenStudy (freckles):

are you asking to write \[\sin(\tan^{-1}(\frac{x}{2})) \text{ as an algebraic expression }\]

OpenStudy (freckles):

the trick is to do: \[u=\tan^{-1}(\frac{x}{2}) \\ \text{ so } \tan(u)=\frac{x}{2} \text{ draw a right triangle and use this information } \] and then find the hyp then to find sin(u), just recall the ratio for it is opp/hyp